Given the equation of a circle as

X^2+Y^2+4x-6y=-


1
Expert's answer
2021-09-10T08:29:52-0400
"x^2+y^2+4x-6y=-4"

"x^2+4x+4+y^2-6y+9=9"

"(x+2)^2+(y-3)^2=9"

The equation of the circle with center "(-2, 3)" and radius "r=3."